Circular No. 8217

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 U 3 A new candidate (R = 25.1–25.3) for an outer satellite of Uranus was

discovered by S. S. Sheppard, University of Hawaii, on images obtained by D. C. Jewitt and himself on Aug. 29 and 30 with the Subaru 8.2-m reflector at Mauna Kea. After they had also observed the object with the Gemini 8.2-m reflector on Sept. 20, B. G. Marsden tentatively identified the object with a satellite candidate observed by M. Holman and J. Kavelaars on the single night of 2001 Aug. 13 with the 4-m Blanco reflector at Cerro Tololo (cf. IAUC 7980, 8213, 8216), a result confirmed by Kavelaars’ recognition and measurement yesterday of a 2001 Aug. 25 image with the Canada-France-Hawaii telescope. Full astrometry, Marsden’s orbital elements (a = 0.098 AU, e = 0.78, i = 51^{o}, H = 12.7), and an ephemeris appear on MPEC 2003-T58.

COMET 157P/TRITTON = P/2003 T1 Total-visual-magnitude and coma-diameter estimates: Oct. 8.17 UT,

11.8, 1′.3 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector; strong central condensation in a faint coma); 8.19, 11.4, 1′.3 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector).

AG DRACONIS J. L. Sokoloski, Harvard-Smithsonian Center for Astrophysics; and

M. P. Rupen and A. J. Mioduszewski, National Radio Astronomy Observatory, on behalf of a larger collaboration, report the detection of the symbiotic star AG Dra in the radio with the Very Large Array (in BnA configuration) on Oct. 7 UT, roughly nine days after the start of an optical outburst. The radio source was not resolved, with a beam size of 0′′.84×0′′.44 FWHM at p.a. 32^{o}. The 8.46-GHz flux density was 0.47 ± 0.09 mJy. The nominal 4.86-GHz flux density was 0.17 ± 0.07 mJy. However, as this 5-GHz flux density was comparable to background fluctuations, it is most conservatively interpreted as a 3σ upper limit of 0.38 mJy. In either case, the 5-GHz flux density has decreased with respect to the value of ≥ 1 mJy found by Ogley et al. (2002, MNRAS 330, 772) in Mar. 2000, when AG Dra was in a quiescent state after a series of optical outbursts. Previous low 5-GHz flux densities were reported by Seaquist and Taylor (1990, Ap.J. 349, 313), when AG Dra was near the peak of a large optical outburst in 1982 (≤ 0.41 mJy) and when it was near the tail end of an optical outburst in 1986 (0.36 ± 0.08 mJy).

2003 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green